Grade 12Mechanics

Two masses m1 and m2 are connected at the two ends of a light intextensible string that passes over a light frictionless fixed pully find the expression for the acceleration of the masses and the tension in the string when the masses are released.

Neha Dhama
9 Years agoGrade 12

1 Answer

Rajessh
9 Years ago
a1=T/m1 and a2=T/m2;
T=m1a1+m1g for downward movement of mass m1
T=m2a2-m2g  for upward movement of mass m2